# Without Paired Labeled Data: An End-to-End Self-Supervised Paradigm for UAV-View Geo-Localization

### v1

@@ III-D Pseudo-Label Enhancement
In an unsupervised learning scenario, the quality of pseudo-labels exerts a decisive influence on model performance. Inspired by [47], this study proposes a pseudo-label optimization framework based on dynamic feature robustness. Specifically, small perturbations are introduced into the original feature representation to form a perturbed feature space, and the Top-K nearest neighbor consistency between the original and perturbed features is employed to identify high-confidence samples while filtering out noisy or less discriminative instances. Subsequently, leveraging the inter-sample similarity matrix in the feature space, a spatial smoothing operation is performed on the initially generated pseudo-labels, thereby enforcing local label consistency among neighboring samples.
We first employe small Gaussian noise $\boldsymbol{\epsilon}_{d},\boldsymbol{\epsilon}_{s}\sim\mathcal{N}(\mathbf{0},\sigma^{2}\mathbf{I})$ into the original satellite instances feature and the original drone instances features to obtain perturbed features, which are used for more robust similarity estimation and subsequent cluster consistency evaluation. specifically:
@@
The label of the $k$-th similar infrared instance in two ranking lists can be represented as:
@@
So, we associate the cross-modality labels with the intersection of two label sets to investigate collaborative ranking consistency:
@@
The reliable refined cross-view label of $\mathbf{f}_{m}^{s}$ can be expressed by the label of maximum count in $I_{\mathbf{f}_{m}^{s}}(k)$:
@@
Then, we convert the label list $\{\tilde{y}_{\mathbf{f}_{m}^{s}}^{cm},m\in[1,M]\}$ to the form of one-hot code matrix $\tilde{Y}^{cm}\in\mathbb{R}^{M\times N}$ by setting the column according to the refined labels to 1 and the rest to 0.Then,two initial and perturbed homogeneous initial $P^{ss}$ and deep $P_{\boldsymbol{\epsilon}}^{ss}$ similarity matrices are constructed to investigate intra-modality ranking consistency, enhancing the precision of refined cross-view pseudo labels by:
@@ IV-A Experimental Datasets and Evaluation Metrics
To evaluate the effectiveness of our proposed cross-view geo-localization framework, we conduct experiments on three benchmark datasets—University-1652 [9], SUES-200 [55], and DenseUAV [53]. These datasets present complementary challenges for multi-view image matching and retrieval.
University-1652 is a large-scale cross-view geo-localization dataset encompassing 1,652 locations from 72 universities worldwide. It contains multi-view images—drone, satellite, and ground-level. Of the total 1,652 buildings, 701 from 33 universities form the training set, while 951 from 39 different universities (with no geographic overlap) comprise the test set. As the first dataset to include drone-view for cross-view geo-localization, it facilitates matching across ground, drone, and satellite perspectives.
SUES-200 centers on altitude diversity in aerial imagery. It includes 200 distinct locations, split into 120 for training and 80 for testing. Each location provides one satellite image and aerial images taken at four altitudes (150m, 200m, 250m, and 300m), covering diverse environments such as urban areas, natural landscapes, and water bodies. SUES-200 thus evaluates cross-view retrieval performance under varying flight heights, reflecting real-world UAV scenarios.
DenseUAV targets UAV self-positioning in low-altitude urban environments, featuring over 27,000 UAV and satellite images from 14 university campuses. It includes dense sampling, multi-scale satellite images, and varying altitudes (80m, 90m, 100m) across multiple temporal conditions. The training set consists of 6,768 UAV images and 13,536 satellite images, while the test set contains 2,331 UAV images and 4,662 satellite images, with no geographic overlap. DenseUAV’s challenges include cross-view matching, neighbor confusion, and spatial-temporal variation, making it a key benchmark for UAV-based localization in GPS-denied settings.
We evaluate the models using Recall@K (R@K) and Average Precision (AP). R@K measures the proportion of correct matches within the Top-$K$ results, while AP balances precision and recall. Model efficiency is assessed through parameter count, reflecting portability under resource constraints. Comparisons are made using the optimal model configurations for each method.
@@ IV-B Implementation Details
We adopt a symmetric sampling strategy for selecting input images. A ConvNeXt-Tiny model, pre-trained on ImageNet, serves as the backbone for feature extraction, with a newly added classifier module initialized via the Kaiming method. During both training and testing, all input images are resized to $3\times 384\times 384$. We employ several data augmentation techniques, including random cropping, random horizontal flipping, and random rotation. The batch size is set to 64, and DBSCAN [10] is used to generate pseudo labels. For optimization, we use the SGD optimizer with an initial learning rate of 0.001, training the model for 30 epochs in total. All experiments are conducted under the PyTorch framework on an Ubuntu 22.04 system equipped with four NVIDIA RTX 4090 GPUs. Additionally, in the University-1652 and SUES-200 datasets, each location provides only a single satellite image; we therefore replicate each satellite feature 50 times for clustering purposes, leaving UAV features unchanged. In DenseUAV, we replicate UAV features four times at each altitude, and replicate the satellite features three times for each time point and altitude.
@@ IV-C Comparison with State-of-the-art Methods
TABLE II: Comparisons between the proposed method and some state-of-the-art methods on the SUES-200 dataset.
@@
[…]
Results on Cross-Domain Generalization Performance: In cross-view geo-localization, cross-domain adaptability is a critical metric for evaluating the generalization capability of a model, especially when the training and testing datasets exhibit significant differences. To evaluate the transferability of our proposed model, we conducted experiments by training on the University-1652 dataset and testing on the SUES-200 dataset. As shown in Table III, our method demonstrates superior generalization performance under the challenging University-1652→SUES-200 and Satellite→Drone cross-domain settings. Remarkably, our model achieves state-of-the-art results without any supervision (0% supervision ratio) while utilizing significantly fewer parameters (28.59M). For University-1652→SUES-200, our method outperforms all fully supervised baselines across most distance thresholds. At 300m, we attain R@1/AP of 96.78%/97.42%, surpassing the previous best results (CAMP) by 1.10% and 0.77%, respectively. Notably, even at lower altitude (150m), our model achieves R@1 of 82.60%, outperforming MEAN and CAMP, highlighting its robustness to domain shifts. In the Satellite→Drone scenario, our method further exhibits exceptional generalization. At 250m, we achieve R@1/AP of 98.75%/94.74%, improving upon the second-best method (DAC) by 3.75% and 1.93%, respectively. The consistent performance gains across varying distances (e.g., 97.50% R@1 at 300m) underscore the effectiveness of our unsupervised paradigm in mitigating domain discrepancies.These results validate that our approach not only reduces dependency on labeled data but also enhances cross-domain adaptability through efficient feature learning. The lightweight architecture (28.59M parameters) further ensures practical applicability in resource-constrained scenarios, setting a new benchmark for unsupervised cross-view geo-localization.
@@
Results on DenseUAV: The DenseUAV dataset presents a highly challenging scenario due to its densely sampled nature, where neighboring images are easily confused, making cross-view matching particularly difficult. Furthermore, the inclusion of multi-scale satellite imagery and UAV images taken at different altitudes adds another layer of complexity, often leading to unstable retrieval performance in conventional methods. However, As shown in Table IV, our approach demonstrates exceptional performance under both fully supervised (100% supervision ratio) and zero-supervision (0% supervision ratio) settings. In the fully supervised case, our method achieves an impressive 99.61% R@1 at 90m altitude and 99.49% R@5 at 100m, closely matching or even surpassing the current state-of-the-art method (MEAN). More notably, under zero supervision, our method still achieves 73.79% R@1 and 90.69% R@5 at All height, significantly outperforming some fully supervised approaches, proving its strong generalization and robustness. These results highlight our method’s ability to effectively address the challenge of neighboring image confusion and provide reliable UAV self-localization in GPS-denied environments.
IV-D Ablation Studies
To comprehensively evaluate the rationale behind our proposed method, we conducted an ablation study on the University-1652 dataset. We systematically analyzed the impact of each proposed module and explored their effectiveness in different configurations to better understand their contributions to the cross-geographic view matching task. The experimental results are shown in Table LABEL:Ablation_Studies.
Backbone: Basic Feature Extraction Capability. The Backbone refers to the performance of the ConvNeXt-Tiny network without training, serving as a direct feature extractor. As shown in Table V, its performance is significantly limited, achieving an R@1 of only 9.55% and 35.24%, and an AP of 11.88% and 13.20% for the Drone $\rightarrow$ Satellite and Satellite $\rightarrow$ Drone tasks, respectively. This result confirms that the backbone alone lacks the capability to extract discriminative features, highlighting the necessity of additional learning mechanisms.
Baseline: Dual-Path Contrastive Learning for Feature Optimization. Incorporating a Baseline model with dual-path cluster-based contrastive learning significantly enhances feature discrimination. The R@1 values rise to 71.06% and 81.31%, while AP improves to 74.91% and 65.26%. These improvements demonstrate that contrastive learning strengthens intra-view feature discrimination but remains insufficient for capturing cross-view consistency.
@@ IV-D Ablation Studies
Dynamic Hierarchical Memory Learning (DHML): Enhancing Global and Local Consistency. To further address consistency between cross-views, we propose DHML. This module effectively mines hierarchical feature dependencies and improves structural alignment. Consequently, R@1 increases to 84.03% and 48.72%, while AP reaches 86.41% and 82.74%. This confirms that DHML significantly strengthens feature robustness across different views.
TABLE V: The influence of each component on the performance of proposed DMNIL.
@@
Information Consistency Evolution Learning (ICEL): Leveraging Cross-View Associations. Although DHML improves feature consistency, it is still constrained to within-view optimizations. Therefore, we propose ICEL to explicitly model cross-view dependencies. ICEL enables mutual feature refinement across views, boosting R@1 to 86.13% and 51.04%, with AP increasing to 88.31% and 82.97%. These results further confirm the effectiveness of ICEL, showing that it significantly enhances information consistency in cross-geographic view matching tasks, improving the overall model performance.
@@
The ablation study comprehensively validates the impact of each component in the proposed DMNIL framework. Starting from the backbone, which lacks discriminative capacity, to the baseline method leveraging contrastive learning, followed by the enhancements brought by DHML for global-local consistency, ICEL for cross-view feature association, and finally PLE for pseudo-label refinement, each module plays a crucial role in improving performance. The results demonstrate that our approach systematically enhances feature learning, leading to superior matching performance on University-1652.
@@
Figure 5: Cluster analysis across different epochs for various datasets. The plot shows the number of clusters for DenseUAV, University-1652, and SUES-200 datasets, each for both drone and satellite data.
@@ IV-E Visualization
Feature space and similarity distribution. To further illustrate the effectiveness of the our method, We perform feature space (t-SNE[56] map) and similarity distribution visualization for DMNIL, as presented in Fig 4.Comparing (a) Baseline and (b) Our Method, it is evident that our approach enhances the clustering of cross-view positive samples, effectively bringing the drone and satellite representations of the same location closer together. Additionally, the separation between different locations is more distinct, indicating better feature discriminability. The second row shows the similarity distribution between drone-satellite positive and negative pairs. Our method demonstrates a clearer distinction between positive and negative pairs, with a reduced overlap between distributions, which suggests improved cross-view matching and reduced feature ambiguity. This highlights the effectiveness of our approach in addressing viewpoint discrepancies across geographic perspectives.
Figure 6: Top-6 Retrieval Results of the Proposed DMNIL on the University-1652 Dataset.
@@
Cluster analysis. To validate the efficacy of our method, as shown in Fig. 5, we illustrate the evolution of cluster counts across epochs for three cross-view datasets. In the DenseUAV dataset (2,255 locations), satellite images exhibit significant clustering fluctuations during the early training stages, due to low intra-class discriminability and weak cross-view correlations. In contrast, the University-1652 dataset (701 locations) demonstrates greater UAV image clustering instability, while satellite image clusters stabilize more rapidly in the initial epochs. The SUES-200 dataset (1,200 locations), with its smaller data size and simpler distribution, exhibits minimal early-stage variation and stabilizes faster than larger datasets. As training progresses, the cluster counts across all datasets gradually converge toward the ground-truth location counts, despite initial inter-class confusion. This process validates our model’s ability to progressively learn discriminative representations.
@@
Retrieval results. To demonstrate the effectiveness of the DMNIL, we present retrieval results on the University-1652 and DenseUAV dataset in Fig. 6 and Fig. 7. Green borders indicate correct matches, while red borders denote errors, showing that DMNIL significantly improves retrieval accuracy.
In the University-1652 dataset, the drone-to-satellite task presents drone queries in the left panel and retrieved satellite images in the right panel. Since each location has only one satellite image, DMNIL consistently achieves Top-1 accuracy, demonstrating its effectiveness. For satellite-to-drone retrieval, DMNIL accurately matches satellite queries to the corresponding drone images (green borders), showcasing its robust cross-view geo-localization capability and strong feature consistency across perspectives.
In the DenseUAV dataset, the retrieval task is more challenging due to dense sampling, multi-scale satellite imagery, and temporal variations, increasing the difficulty of cross-view matching. Since each location corresponds to multiple satellite images, the model must handle neighbor confusion effectively. In the Drone $\rightarrow$ Satellite task, despite variations in UAV altitude and satellite scale, our method accurately extracts spatial-geographic features, achieving high retrieval accuracy and significantly reducing mismatches. Compared to University-1652, DenseUAV better reflects real-world UAV localization scenarios, where our method demonstrates strong cross-view retrieval performance, effectively distinguishing dense regions and enhancing model robustness.
Figure 7: Top-6 Retrieval Results of the Proposed DMNIL on the DenseUAV Dataset.

### v2

@@ III-D Pseudo-Label Enhancement
In self-supervised learning, the quality of pseudo-labels exerts a decisive influence on model performance. Inspired by [52], our proposes a pseudo-label optimization strategy based on dynamic feature robustness. Specifically, we introduce small perturbations to the original feature representations and exploit the top-K neighbor consistency between the original and perturbed features to identify high-confidence samples while filtering out noisy instances. Then, a spatial smoothing operation based on the similarity matrix is applied to refine the pseudo-labels, enforcing local label consistency.
We first employe small Gaussian noise $\boldsymbol{\epsilon}_{d},\boldsymbol{\epsilon}_{s}\sim\mathcal{N}(\mathbf{0},\sigma^{2}\mathbf{I})$ into the original satellite instances feature and the original drone instances features to obtain perturbed features, which are used for more robust similarity estimation and subsequent cluster consistency evaluation. It can be formulated as
@@
The label of the $k$-th similar infrared instance in two ranking lists can be represented as
@@
So, we associate the cross-modality labels with the intersection of two label sets to investigate collaborative ranking consistency. This process can be described as
@@
The reliable refined cross-view label of $\mathbf{f}_{m}^{s}$ can be expressed by the label of maximum count in $I_{\mathbf{f}_{m}^{s}}(k)$.
@@
Then, we convert the label list $\{\tilde{y}_{\mathbf{f}_{m}^{s}}^{cm},m\in[1,M]\}$ to the form of one-hot code matrix $\tilde{Y}^{cm}\in\mathbb{R}^{M\times N}$ by setting the column according to the refined labels to 1 and the rest to 0.Then,two initial and perturbed homogeneous initial $P^{ss}$ and deep $P_{\boldsymbol{\epsilon}}^{ss}$ similarity matrices are constructed to investigate intra-view ranking consistency, enhancing the precision of refined cross-view pseudo labels by:
@@ IV-A Experimental Datasets and Evaluation Metrics
To evaluate our self-supervised method, we conduct experiments on three benchmark datasets including University-1652 [12], SUES-200 [33] and DenseUAV [54].
University-1652 contains images from 1,652 buildings across 72 universities worldwide. It includes drone, satellite, and ground-level views. The training set consists of 701 buildings from 33 universities, and the test set covers 951 buildings from 39 geographically disjoint universities.
SUES-200 focuses on altitude diversity in aerial imagery, containing 200 locations with 120 for training and 80 for testing. Each location provides one satellite image and aerial images captured at four altitudes (150m, 200m, 250m, and 300m), covering urban areas, natural landscapes, and water bodies. SUES-200 evaluates cross-view retrieval under varying flight heights, reflecting real-world UAV scenarios.
DenseUAV focuses on UAV self-positioning in low-altitude urban environments, containing over 27,000 UAV and satellite images collected from 14 university campuses. It features dense sampling, multi-scale satellite images, and varying altitudes (80m, 90m, and 100m) under multiple temporal conditions. The training set includes 6,768 UAV images and 13,536 satellite images, while the test set comprises 2,331 UAV images and 4,662 satellite images from non-overlapping areas. DenseUAV poses challenges such as cross-view matching, neighbor confusion, and spatial-temporal variation.
We evaluate the models using Recall@K (R@K) and Average Precision (AP), where R@K measures the proportion of correct matches among the Top-$K$results and AP balances precision and recall. Model efficiency is assessed by the number of parameters, reflecting portability under resource constraints. All comparisons are conducted using the best configurations for each method.
@@ IV-B Implementation Details
A ConvNeXt-Tiny model, pre-trained on ImageNet, serves as the backbone for feature extraction, with a newly added classifier module initialized via the Kaiming method. During both training and testing, all input images are resized to $3\times 384\times 384$. We employ several data augmentation techniques, including random cropping, random horizontal flipping, and random rotation. The batch size is set to 64, and DBSCAN [8] is used to generate pseudo labels. For optimization, we use the SGD optimizer with an initial learning rate of 0.001, training the model for 30 epochs in total. All experiments are conducted under the PyTorch framework on an Ubuntu 22.04 system equipped with four NVIDIA RTX 4090 GPUs. Additionally, in the University-1652 and SUES-200 datasets,we replicate each satellite feature 50 times for clustering purposes, leaving UAV features unchanged. In DenseUAV, we replicate UAV features four times at each altitude, and replicate the satellite features three times for each time point and altitude.
@@ IV-C Comparison with State-of-the-art Methods
To evaluate the effectiveness of our method, we compare it with the state-of-the-art supervised methods MuSe-Net [54], MCCG [1], SDPL [3], CCR [32], Sample4Geo [29], SRLN[31], MEAN [2] and DAC [4], and the recent self-supervised methods EM-CVGL [5], CDIKTNet [42].
@@
We further benchmark against eight representative supervised methods, and the results show that our method clearly surpasses some of them, including MuSe-Net, MCCG, and SDPL. Furthermore, with self-supervised pre-training and subsequent supervised fine-tuning, our method achieves competitive performance under the supervised setting, comparable to CCR, Sample4Geo, and SRLN.
[…]
@@
Results on SUES-200. As shown in Table II, As shown in Table II, our method, when directly transferring the pre-trained model from the University-1652 dataset to the SUES-200 dataset without any re-training, consistently outperforms all existing self-supervised methods and achieves competitive performance compared to supervised methods such as MCCG and SDPL. Moreover, by performing self-supervised re-training on the SUES-200 dataset, our method brings a significant performance boost, achieving results comparable to the supervised CCR. Notably, when further fine-tuning the transferred model for only one epoch, we observe additional improvements, surpassing the supervised SRLN. These results clearly demonstrate the effectiveness of our method in self-supervised learning, transfer learning, and efficient fine-tuning.
Results on Cross-Domain Generalization Performance: Our method consistently outperforms existing self-supervised method under cross-domain settings, showing both significant improvements and stable generalization. In the Drone to Satellite setting, our model achieves better performance than EM-CVGL at all altitudes. At 150 meters, R@1 and AP increase by approximately 22% and 20%; at 200 meters, the gains are around 19% and 16%; at 250 meters, nearly 14% and 12%; and even at 300 meters, our method still outperforms by about 11% and 9%. In the Satellite to Drone scenario, similar improvements are observed. At 150 meters, R@1 and AP improve by approximately 18% and 25%; at 200 meters, by 10% and 20%; at 250 meters, by 6% and 13%; and at 300 meters, by 3% and 10%. These results confirm the robustness and strong cross-domain generalization capability of our method across varying altitudes.Remarkably, our method even surpasses all existing state-of-the-art supervised methods in terms of cross-domain generalization. These results confirm the robustness and strong cross-domain generalization capability of our method across varying altitudes.
@@
Results on DenseUAV: The DenseUAV dataset presents a highly challenging scenario due to its densely sampled nature, where neighboring images exhibit high visual similarity, often leading to severe confusion in UVGL. In addition, the dataset includes satellite images captured at different times and UAV images from multiple altitudes, which further increases intra-domain variability and task complexity. As shown in Table IV, under the fully unsupervised setting, our method achieves 73.79% R@1 and 90.69% R@5 across all altitudes, significantly outperforming existing self-supervised methods. It is noteworthy that the two representative self-supervised baselines consistently yield poor performance on DenseUAV, highlighting their limitations in handling densely sampled and spatiotemporal diverse scenarios. Furthermore, after supervised fine-tuning, our method surpasses several supervised baselines, including ConvNeXt-Tiny, Sample4Geo, CAMP, and DAC. In some cases, it even outperforms the current best-performing supervised method, MEAN. These results clearly demonstrate the effectiveness of our method in addressing extreme feature confusion caused by spatial continuity and appearance shifts induced by temporal variation, confirming its strong generalization and robustness.
IV-D Ablation Studies
To thoroughly evaluate the effectiveness of our proposed method, we conduct an ablation study on the University-1652 dataset. We systematically assess the impact of each individual module and investigate their contributions under different configurations. The results are presented in Table LABEL:Ablation_Studies.
ConvNeXt-Tiny Backbone. The Backbone refers to the untrained ConvNeXt-Tiny network used as a direct feature extractor. As shown in Table LABEL:Ablation_Studies, its performance is significantly limited, achieving an R@1 of only 9.55% and 35.24%, and an AP of 11.88% and 13.20% for the Drone $\rightarrow$ Satellite and Satellite $\rightarrow$ Drone, respectively. These results indicate that the backbone alone lacks the ability to extract discriminative cross-view features.
Effectiveness of Dual-Path Contrastive Learning Baseline. After incorporating the dual-path contrastive learning baseline, the R@1 improves to 71.06% and 81.31%, while AP reaches 74.91% and 65.26%, respectively. Although the baseline primarily learns intra-view feature discrimination, the training process also facilitates a certain degree of alignment between UAV and satellite representations. This leads to the formation of a more consistent and unified feature representation space, which serves as a solid foundation for subsequent cross-view learning.
Effectiveness of Dynamic Hierarchical Memory Learning. Compared with the contrastive learning baseline, incorporating DHML leads to significant performance improvements under both settings. Specifically, R@1 increases by 12.97% and 9.13%, and AP improves by 11.5% and 17.48% , reaching 84.03% and 90.44% in R@1, and 86.41% and 82.74% in AP, respectively. These gains can be attributed to the joint modeling of long-term and short-term feature dependencies introduced by DHML, which enhances the capacity to capture spatial structures in UVGL scenarios and enhances intra-view feature representation and discrimination.
TABLE V: The influence of each component on the performance of proposed method.
@@ IV-D Ablation Studies
Effectiveness of Information Consistency Evolution Learning. Although DHML enhances intra-view feature representation and discrimination, its optimization remains limited to within-view information. To address this limitation, ICEL is proposed to learn cross-view dependencies. By enabling bidirectional information interaction between different views, ICEL facilitates cross-view feature refinement. As a result, R@1 improves to 86.13% and 51.04%, while AP reaches 88.31% and 82.97%. These results further validate the effectiveness of ICEL in enhancing cross-view consistency for UVGL, thereby contributing to overall performance improvement.
@@
The ablation study systematically evaluates the contribution of each component in the proposed self-supervised framework. Starting with a backbone that lacks sufficient discriminative power, we first establish a dual-branch contrastive learning baseline. On top of this, DHML is introduced to strengthen intra-view feature representation and discriminability, ICEL is employed to facilitate cross-view feature association, and PLE is leveraged for pseudo-label refinement. Experimental results show that our method significantly enhances the learning of discriminative and view-consistent representations under self-supervised learning.
@@ IV-E Visualization
Feature space and similarity distribution. To further illustrate the effectiveness of the our method, We perform feature space (t-SNE[55] map) and similarity distribution visualization for our method, as presented in Fig 4.Comparing (a) Baseline and (b) Ours, it is evident that our method enhances the clustering of cross-view positive samples, effectively bringing the drone and satellite representations of the same location closer together. Additionally, the separation between different locations is more distinct, indicating better feature discriminability. The second row shows the similarity distribution between drone-satellite positive and negative pairs. Our method demonstrates a clearer distinction between positive and negative pairs, with a reduced overlap between distributions, which suggests improved cross-view matching and reduced feature ambiguity. This highlights the effectiveness of our method in UVGL.
@@
Cluster analysis. To validate the efficacy of our method, as shown in Fig. 5, we illustrate the evolution of cluster counts across epochs for three cross-view datasets. In the DenseUAV dataset (2,255 locations), satellite images exhibit significant clustering fluctuations during the early training stages, due to low intra-class discriminability and weak cross-view correlations. In contrast, the University-1652 dataset (701 locations) demonstrates greater UAV image clustering instability, while satellite image clusters stabilize more rapidly in the initial epochs. The SUES-200 dataset (1,200 locations), with its smaller data size and simpler distribution, exhibits minimal early-stage variation and stabilizes faster than larger datasets. As training progresses, the cluster counts across all datasets gradually converge toward the ground-truth location counts, despite initial inter-class confusion. This process validates the ability of our method to progressively learn discriminative representations.
Figure 5: Cluster analysis across different epochs for various datasets. The plot shows the number of clusters for DenseUAV, University-1652, and SUES-200 datasets.
Figure 6: Top-6 retrieval results of the proposed method on the University-1652 dataset.
Figure 7: Top-6 retrieval results of the proposed method on the DenseUAV dataset.
@@
Retrieval results. To demonstrate the effectiveness of our method, we present retrieval results on the University-1652 and DenseUAV dataset in Fig. 6 and Fig. 7. Green borders indicate correct matches, while red borders denote errors, showing that our method significantly improves retrieval accuracy.
[…]
In the DenseUAV dataset, the retrieval task becomes significantly more challenging due to dense spatial sampling, multi-scale satellite imagery, and temporal dynamics, all of which exacerbate the difficulty of cross-view representation learning. In the Drone $\rightarrow$ Satellite retrieval setting, despite variations in UAV altitudes and satellite image resolutions, our method effectively captures spatial-geographic cues, achieving high retrieval accuracy while substantially reducing mismatches. Compared to University-1652, DenseUAV provides a more realistic and complex benchmark for UAV-based geo-localization. Under this setting, our method exhibits strong cross-view retrieval performance, demonstrating robust feature discrimination in densely distributed urban regions.